# Hydra: Breaking the Global Ordering Barrier in Multi-BFT Consensus

### v1

@@ IX Related Work
Multi-BFT consensus. The foundation of Byzantine fault-tolerant consensus was established by Castro and Liskov in PBFT [pbft1999], which inspired a long line of practical protocols such as Zyzzyva [kotla2007zyzzyva], Tendermint [Buchman2016TendermintBF], and HotStuff [hotstuff]. These leader-based protocols streamline the process of reaching agreement but inevitably concentrate decision power on a single replica, resulting in a leader bottleneck that constrains throughput and scalability [alqahtani2021bottlenecks, charapko2021pigpaxos, gai2023scaling, kang2025hotstuff, gupta2023dissecting].
@@
To alleviate this limitation, Multi-BFT consensus architectures were introduced, allowing replicas to run several consensus instances concurrently and thereby removing the dependence on a single leader [stathakopoulou2022state, gupta2021rcc, MIR-BFT, Ladon2025, kang2024spotless]. Among early designs, Mir-BFT [MIR-BFT] pioneered this direction by executing multiple instances in parallel and establishing a global order of blocks based on pre-assigned indices. However, because Mir-BFT triggers an epoch change whenever any instance leader behaves incorrectly, the system remains vulnerable to Byzantine disruptions. ISS [stathakopoulou2022state] refined this model by introducing no-op deliveries to mitigate unnecessary epoch changes, while RCC [gupta2021rcc] similarly relies on a static global order of outputs. In both cases, a slow or faulty instance can delay epoch progression for all replicas, creating performance bottlenecks.
@@
Other variants attempt to optimize the global ordering process. For example, DQBFT [dqbft] dedicates a separate BFT instance to globally order outputs from other instances, simplifying coordination but making the system more prone to targeted attacks [estrada2006network]. Ladon [Ladon2025] introduces monotonic ranking to dynamically determine cross-instance ordering, mitigating straggler-induced delays. Building upon this, Orthrus [Orthrus2025] further introduces a fast path for independent transactions, enabling them to bypass the global ordering phase, while still maintaining global consistency for dependent ones. Although these techniques alleviate synchronization overheads, the global ordering phase still constitutes a large fraction of overall system latency.
@@
Unlike the systems above, TELL [tong2024tell] shifts optimization efforts to the execution layer instead of the consensus protocol itself. It employs a State Hash Table (SHT) to track read/write dependencies, enabling concurrent execution both across instances and within blocks. Conflicts are handled by selective re-execution and periodic merging of instance states at the epoch boundary. While this strategy reduces execution wait time, its improvement in end-to-end latency remains limited due to the overhead of reprocessing conflicting transactions.
@@
Partial ordering design. A complementary research direction focuses on partially ordered execution to relax the global ordering requirement, particularly within payment-oriented systems. CryptoConcurrency [tonkikh2023cryptoconcurrency] dynamically detects overspending by checking account balances, enabling concurrent transactions without full consensus. Astro [collins2020online] maintains per-client logs to prevent double spending, while ABC [sliwinski2021asynchronous] allows validators to process transactions in parallel without global coordination, thereby enhancing efficiency. FastPay [baudet2020fastpay] leverages payment semantics to minimize shared state between accounts, facilitating asynchronous execution with high concurrency. Flash [lewis2023flash] bypasses reliable broadcast through a DAG-based, partially ordered structure. A non-sequential model for monetary transfers is proposed in [auvolat2020money], based on reliable broadcast abstraction. Pastro [kuznetsov2023permissionless] further defines a partially ordered transaction set that determines active participants and stake distribution, offering flexibility for applications that do not depend on a total order. While these approaches achieve efficient execution for payment systems, they cannot generally support complex smart contract semantics.
@@
Hybrid ordering design. Recent advances in blockchain systems have explored hybrid ordering mechanisms that selectively apply global ordering only when necessary [blackshear2023sui, babel2025mysticeti]. Sui Lutris [blackshear2023sui] introduces a dual-path model, where single-owner transactions follow a lightweight fast path, while shared-object transactions are ordered through a consensus path. This design reduces latency for independent operations but depends on client-side orchestration, which requires gathering additional certificate signatures and performing periodic checkpointing to ensure finality. Mysticeti [babel2025mysticeti] generalizes the same concept using a DAG-based consensus protocol, offering a tighter integration between the fast and consensus paths. However, Mysticeti still separates execution and finality: locally executed transactions may later be reverted if not sufficiently confirmed at epoch boundaries, complicating persistence across epochs. Orthrus [Orthrus2025] extends this hybrid ordering paradigm into a Multi-BFT framework while retaining a standard BFT core. It introduces a fast path where conflict-free transactions can be finalized immediately upon instance commitment, thus eliminating deferred confirmation and preventing rollback between epochs. Thunderbolt [chen2024thunderbolt] introduces a hybrid ordering mechanism that deterministically orders cross-shard transactions while preserving partial concurrency for single-shard execution, achieving both high scalability and consistency across shards.

### v2

@@ IX Related Work
Multi-BFT consensus. The foundation of Byzantine fault-tolerant consensus was established by Castro and Liskov in PBFT [14], which inspired a long line of practical protocols such as Zyzzyva [38], Tendermint [12], and HotStuff [59]. These leader-based protocols streamline the process of reaching agreement but inevitably concentrate decision power on a single replica, resulting in a leader bottleneck that constrains throughput and scalability [2, 16, 23, 34, 30].
@@
To alleviate this limitation, Multi-BFT consensus architectures were introduced, allowing replicas to run several consensus instances concurrently and thereby removing the dependence on a single leader [52, 29, 51, 43, 35]. Among early designs, Mir-BFT [51] pioneered this direction by executing multiple instances in parallel and establishing a global order of blocks based on pre-assigned indices. However, because Mir-BFT triggers an epoch change whenever any instance leader behaves incorrectly, the system remains vulnerable to Byzantine disruptions. ISS [52] refined this model by introducing no-op deliveries to mitigate unnecessary epoch changes, while RCC [29] similarly relies on a static global order of outputs. In both cases, a slow or faulty instance can delay epoch progression for all replicas, creating performance bottlenecks.
@@
Other variants attempt to optimize the global ordering process. For example, DQBFT [4] dedicates a separate BFT instance to globally order outputs from other instances, simplifying coordination but making the system more prone to targeted attacks [20]. Ladon [43] introduces monotonic ranking to dynamically determine cross-instance ordering, mitigating straggler-induced delays. Building upon this, Orthrus [42] further introduces a fast path for independent transactions, enabling them to bypass the global ordering phase, while still maintaining global consistency for dependent ones. Although these techniques alleviate synchronization overheads, the global ordering phase still constitutes a large fraction of overall system latency.
@@
Unlike the systems above, TELL [53] shifts optimization efforts to the execution layer instead of the consensus protocol itself. It employs a State Hash Table (SHT) to track read/write dependencies, enabling concurrent execution both across instances and within blocks. Conflicts are handled by selective re-execution and periodic merging of instance states at the epoch boundary. While this strategy reduces execution wait time, its improvement in end-to-end latency remains limited due to the overhead of reprocessing conflicting transactions.
@@
Partial ordering design. A complementary research direction focuses on partially ordered execution to relax the global ordering requirement, particularly within payment-oriented systems. CryptoConcurrency [54] dynamically detects overspending by checking account balances, enabling concurrent transactions without full consensus. Astro [18] maintains per-client logs to prevent double spending, while ABC [49] allows validators to process transactions in parallel without global coordination, thereby enhancing efficiency. FastPay [8] leverages payment semantics to minimize shared state between accounts, facilitating asynchronous execution with high concurrency. Flash [40] bypasses reliable broadcast through a DAG-based, partially ordered structure. A non-sequential model for monetary transfers is proposed in [5], based on reliable broadcast abstraction. Pastro [39] further defines a partially ordered transaction set that determines active participants and stake distribution, offering flexibility for applications that do not depend on a total order. While these approaches achieve efficient execution for payment systems, they cannot generally support complex smart contract semantics.
@@
Hybrid ordering design. Recent advances in blockchain systems have explored hybrid ordering mechanisms that selectively apply global ordering only when necessary [11, 7]. Sui Lutris [11] introduces a dual-path model, where single-owner transactions follow a lightweight fast path, while shared-object transactions are ordered through a consensus path. This design reduces latency for independent operations but depends on client-side orchestration, which requires gathering additional certificate signatures and performing periodic checkpointing to ensure finality. Mysticeti [7] generalizes the same concept using a DAG-based consensus protocol, offering a tighter integration between the fast and consensus paths. However, Mysticeti still separates execution and finality: locally executed transactions may later be reverted if not sufficiently confirmed at epoch boundaries, complicating persistence across epochs. Orthrus [42] extends this hybrid ordering paradigm into a Multi-BFT framework while retaining a standard BFT core. It introduces a fast path where conflict-free transactions can be finalized immediately upon instance commitment, thus eliminating deferred confirmation and preventing rollback between epochs. Thunderbolt [17] introduces a hybrid ordering mechanism that deterministically orders cross-shard transactions while preserving partial concurrency for single-shard execution, achieving both high scalability and consistency across shards.
@@ XI AI-Generated Content Acknowledgment
Appendix A Correctness Analysis
Lemma 1.
Proof.
Lemma 2.
Proof.
Case 2: Deadlock Resolution. In this case, $tx$ is aborted due to a deadlock. By Lemma 1, all honest replicas will compute the same deadlock group $D$ by invoking $\mathsf{expandDeadlockGroup}$(txtx). All honest replicas apply the same deterministic victim selection rule (e.g., selecting the lowest-indexed transactions) in $\mathsf{selectVictims}$($D$). Since both $D$ and the selection rule are consistent across replicas, the resulting victim set is identical at all honest replicas. Therefore, if $tx\in\mathsf{selectVictims}(D)$ at one honest replica, then all replicas will also abort $tx$.
Theorem 1 (Safety).
Proof.
Lemma 3.
Proof.
Lemma 4.
If a transaction $tx$ is confirmed, it will eventually be executed.
Proof.
Theorem 2 (Liveness).
Proof.
Appendix B Additional Experiments
(a) Throughput
(b) Latency
Figure 8: Throughput and latency of Hydra with increasing cross-instance complexity.
Cross-instance transactions spanning more instances. We evaluate scenarios in which each cross-instance transaction spans 2–5 instances, representing increasingly complex coordination patterns.
As shown in Fig. 8, transactions spanning more instances lead to a moderate reduction in throughput and an increase in latency compared to transactions involving fewer instances, but the degradation remains bounded. Specifically, as the number of involved instances increases from 2 to 5, throughput decreases approximately 11.8%. Over the same range, latency increases about 13.9%.
These results indicate that Hydra can efficiently support cross-instance transactions involving a larger number of instances, and that increasing cross-instance transaction complexity does not introduce significant additional performance degradation.
Appendix C Comparison with Sharding
This appendix clarifies the relationship between Hydra and sharding designs.
C-A Cross-Shard Transaction Handling
Existing sharding designs fundamentally rely on coordination among the involved shards to guarantee atomicity and consistency of cross‑shard transactions. We next discuss several representative works along these directions.
Client‑driven two‑phase commit (2PC). This class originates from systems such as OmniLedger [37]. In these designs, the client orchestrates a two‑phase commit across shards. In the prepare phase, the client sends the transaction to all input shards; after each shard verifies the transaction and locks the necessary state, it returns a signed proof. In the commit phase, the client aggregates these proofs and sends a combined proof to the output shards to finalize the transaction.
Committee‑driven two‑phase commit. In this category, shards coordinate among themselves without relying on clients for orchestration [1, 31, 41]. For example, in Chainspace, where committees of replicas in each shard perform BFT consensus locally and then exchange signed messages among committees to drive prepare and commit decisions. Although this removes client dependency, it still requires explicit cross‑shard communication and multiple coordination rounds.
Intersection replicas. Some protocols introduce structural overlap in the validator sets of shards. Approaches such as Pyramid [32] select replicas that participate in multiple shards simultaneously, enabling them to act as natural bridges for cross‑shard transactions. These intersection replicas can validate and confirm cross‑shard state changes within a single consensus epoch, reducing the need for explicit message exchange across shards. This technique trades off increased validator complexity and careful committee assignment for lower coordination overhead.
Deterministic ordering. Several recent high‑performance blockchain designs depart from traditional sharded BFT architectures by avoiding explicit network or committee partitioning, and instead rely on deterministic ordering to enable parallel execution [24, 11]. As a result, these designs are conceptually closer to the Multi-BFT paradigm. In these systems, transactions that may conflict are assigned a pre-determined global order. By eliminating dynamic coordination during execution, this approach substantially reduces the number of communication rounds required. However, it still incurs overhead for establishing and distributing the global order.
Table IV: Cross-shard / cross-instance transaction coordination overhead in representative BFT systems. $n$: number of replicas, $k$: number of shards/instances.
System Coordination messages
OmniLedger [37] $O(k)$
Chainspace [1] $O(k^{2})$
Pyramid [32] $O(n)$
Hydra nonenone
C-B Coordination Overhead Comparison
To compare different designs, we consider a transaction that spans $k$ shards/instances and estimate the number of extra coordination messages of different protocols in Table IV.
  • In OmniLedger [37], the client collects Proof-of-Acceptance from all input shards and submits the transaction to the output shard. Each involved shard lerader sends at least one message to the client, resulting in a total coordination message complexity of $O(k)$.
  • During the prepare phase of Chainsapce [1], each shard leader needs to validate with the other k1k-1 shards, and then subsequently submits the collected lock proofs to the output shards. In the worst case, the total message complexity is $O(k^{2})$, which comes from shard-to-shard verification.
  • Pyramid [32] selects intersection replicas that belong to multiple shards to process cross-shard transactions. Each intersection node participates simultaneously in the BFT consensus of the relevant shards,and then synchronize the cross-shard commit within each shard. This yields $O((n/k)\cdot k)=O(n)$ messages, roughly linear in total system size.
  • Hydra does not require any cross-instance coordination messages. Cross-instance transactions are executed correctly using locks and deadlock-resolution mechanisms without sending additional messages. Therefore, the coordination messages can be considered as none.